OpenStudy (alante):

What is the solution of the system? 12x - 3y = 18 8x = 2y + 12 A. No solution B. Infinitely many solutions C. (4, 3) D. (3, 4)

OpenStudy (jjcoolja12):

Problem: Solve 12x−3y=18;8x=2y+12 Steps: I will try to solve your system of equations. 12x−3y=18;8x=2y+12 Step: Solve12x−3y=18for x: 12x−3y+3y=18+3y(Add 3y to both sides) 12x=3y+18 12x 12 = 3y+18 12 (Divide both sides by 12) x= 1 4 y+ 3 2 Step: Substitute 1 4 y+ 3 2 forxin8x=2y+12: 8x=2y+12 8( 1 4 y+ 3 2 )=2y+12 2y+12=2y+12(Simplify both sides of the equation) 2y+12+−2y=2y+12+−2y(Add -2y to both sides) 12=12 12+−12=12+−12(Add -12 to both sides) 0=0 Answer: Infinitely many solutions.

OpenStudy (jjcoolja12):

so its b